# GOALS-JWST: The Warm Molecular Outflows of the Merging Starburst Galaxy NGC 3256

### v1

@@ Abstract
We present James Webb Space Telescope (JWST) Integral Field Spectrograph observations of NGC 3256, a local infrared-luminous late-stage merging system with two nuclei about 1 kpc apart, both of which have evidence of cold molecular outflows. Using JWST NIRSpec and MIRI datasets, we investigate this morphologically complex system on spatial scales of 100 pc, where we focus on the warm molecular H gas surrounding the nuclei. We detect collimated outflowing warm H gas originating from the southern nucleus, though we do not find significant outflowing warm H gas surrounding the northern nucleus. Within the observed region, the maximum intrinsic velocities of the outflow reach up to 1,000 km s, and extend out to a distance of 0.7 kpc. Based on H S(7)/S(1) line ratios, we find a larger fraction of warmer gas near the S nucleus, which decreases with increasing distance from the nucleus, signifying the S nucleus as a primary source of H heating. The gas mass of the warm H outflow component is estimated to be = 8.910, as much as 4 of the cold H mass as estimated using ALMA CO data. The outflow time scale is about yr, resulting in a mass outflow rate of = 1.3 M yr and kinetic power of erg s. Lastly, the regions where the outflowing gas reside show high [Fe ii]/Pa and H/Br line ratios, indicating enhanced mechanical heating caused by the outflows. At the same time, the 3.3 m and 6.2 m Polycyclic Aromatic Hydrocarbon fluxes in these regions are not significantly suppressed compared to those outside the outflows, suggesting the outflows have no clear negative feedback effect on the local star formation.
@@ 1 Introduction
The physical processes that govern the evolution of galaxies are of fundamental importance to our understanding of how galaxies in the Universe formed. These processes manifest in many forms, including internal mechanisms, such as stellar and AGN outflows (e.g., Müller-Sánchez et al., 2011; U et al., 2019; Aravindan et al., 2023), and external processes such as galaxy mergers (e.g., Lin et al., 2008; Privon et al., 2013; Conselice, 2014). Outflows are believed to regulate and suppress star formation and can, in many instances, lead their hosts to the well-defined red sequence (e.g., Croton et al., 2006; King & Pounds, 2015; Veilleux et al., 2020). Major mergers also play a critical role in the formation of massive galaxies, and are attributed to the co-evolution of the host galaxy and supermassive black holes (e.g., Medling et al., 2015; Ricci et al., 2017; Ramos Almeida et al., 2019) and the formation of compact star clusters (e.g., Mulia et al., 2015; Adamo et al., 2020; Linden et al., 2021). Local luminous infrared galaxies (LIRGs, ) are ideal systems for studying these processes because they allow us to analyze evolutionary processes across the full array of merger states, and evaluate the elevated star formation rates triggered by such mergers. Their diversity in starburst and AGN power, merger stages, and outflow activity makes LIRGs compelling ecosystems for tying these properties to the stages of galaxy evolution.
@@
One of the most compelling aspects of NGC 3256 is the presence of outflows in both nuclei. Submillimeter ALMA observations indicate two different biconical outflows of molecular CO gas originating from the nuclei, with deprojected maximum velocities (and inclination angles) of 750 (60) and 2600 km s (20) for the outflows in the northern and southern galaxies, respectively (Sakamoto et al., 2014; Michiyama et al., 2018). Near-infrared (NIR) VLT/SINFONI observations of the S nucleus also reveal biconical outflows in the warm, 2000 K H gas, where the gas mass of these outflowing components are estimated to be 1,200 (Emonts et al., 2014).
While previous observations have been able to analyze the warm and cold molecular gas phases in NIR and submillimeter wavelengths with great detail, analysis of the warm H component in the mid-infrared (MIR) has been mainly restricted to using the S(1)–S(3) transitions due to the resolution and sensitivity of the Spitzer Space Telescope (e.g. Stierwalt et al., 2013; Inami et al., 2013). Petric et al. (2018) calculated the mass of the warm molecular H component of the northern galaxy to be log() = 7.38 and a temperature of 322 K within the IRS slits that were centered on the N nucleus. While Spitzer gave a great first look at the warm H gas, the H emission was not spatially resolved and the spectral resolution of Spitzer, R60–600, made it difficult to spectrally resolve the outflowing gas in NGC 3256.
The various H transitions and their line ratios with other IR lines are useful for identifying the sources heating the molecular gas and energizing the ISM (Larkin et al., 1998; Riffel et al., 2013; Colina et al., 2015). They can be used to identify the presence of shocks induced by the outflows, and help evaluate the role outflows have in regulating star formation. With JWST, we can now perform the most detailed, spatially resolved analysis of outflowing warm H gas to date in NGC 3256. The unprecedented spatial resolution of JWST enables us to examine the warm H component on scales of 40–100 pc. We also take advantage of the superb sensitivity and wavelength coverage of JWST to analyze the full set of H 0–0 rotational lines between 5–17 m. Specifically, we utilize integral-field spectroscopy (IFS) of the Near-InfraRed Spectrograph (NIRSpec, Böker et al., 2022; Jakobsen et al., 2022) and medium-resolution spectroscopy (MRS) of the Mid-Infrared Instrument (MIRI, Rieke et al., 2015; Labiano et al., 2021; Wright et al., 2023) to assess the kinematics and energetics of the outflows, and evaluate their impact on the local interstellar medium (ISM).
In our analysis, we adopt a cosmology of km s Mpc, , and . At the redshift of NGC 3256, (40.4 Mpc), subtends 190 pc.
@@ 2 Observations and Data Reduction
Observations of NGC 3256 were taken as part of the ERS Program 1328 (Co-PIs: L. Armus and A. Evans), for which the JWST/NIRSpec IFS data were taken on 2022 December 23-24 UT. Two sets of observations, each centered on a nucleus, were taken using three high-resolution gratings: G140H/F100LP, G235H/F170LP, and G395H/F290LP. This resulted in a wavelength coverage of 0.97–5.27 m, with a nominal resolving power of R2700. A 4-point dither pattern was utilized, and total exposure times for all filters were 934 s and 2218 s for the northern and southern pointings, respectively. Observations using JWST/MIRI in the MRS mode were taken on 2022 December 27 UT. Like for NIRSpec, a separate pointing for each nucleus was taken. These observations include all three sub-bands (short [A], medium [B], and long [C]) in all four channels, which cover the full wavelength range from 4.9–28.8 m. The total exposure time for each sub-band/channel was 444 s, and was the same for both pointings. In this article, we predominately use channels 1–3, which have a nominal resolving power of R2000–3700.
The data were processed through the standard JWST Calibration Pipeline (ver.1.12.5 for NIRSpec; ver.1.11.3 for MIRI, Bushouse et al., 2022). Stage 1 of the reduction pipeline, Detector1, implements detector-level corrections and outputs count-rate data products. Stage 2 (spec2) applies various calibrations steps, including distortion, wavelength, and flux calibrations. Stage 3 (spec3) performs background subtraction for MIRI data and builds the final data cubes from the individual exposures. We also incorporated leakcal calibration in our NIRSpec data. Before stage 3 was implemented, we only included spaxels with data quality (DQ) equal to 0 in order to remove bad spaxels contaminating the NIRSpec data products. We additionally flagged any spaxels in the post-pipeline products with values significantly above emission lines or that were negative.
To properly align both the NIRSpec and MIRI data, Gaia DR3 (Gaia Collaboration et al., 2016, 2023) reference frames of nearby field stars were first used to derive WCS corrections for the JWST/NIRCam and MIRI images (also obtained by ERS 1328, see Linden et al., in prep.). The IFS cubes were subsequently collapsed into a single frame and aligned to the WCS corrected NIRCam/MIRI images, where corrections were around RA: + and DEC: + for NIRSpec, and RA: and DEC: + for MIRI.
Lastly, fringing corrections were applied to all MIRI spectra after extraction. We also stitched the spectra by scaling all the continua to that of MIRI channel 3 long, the longest wavelength channel used in our analysis, where scaling factors were typically around 5.
@@ 3 Methods of Analysis
The primary focus of this article is to obtain robust mass estimates and temperatures of the warm outflowing H gas, as well as analyze its kinematics and energetics. Due to the complex kinematic nature of this merging system, robust emission line decompositions and detailed mapping of the spatial structures are needed to accurately identify and study the H gas. To accomplish this, we used the Bayesian AGN Decomposition Analysis for SDSS Spectra (BADASS111https://github.com/remingtonsexton/BADASS3, Sexton et al., 2021) code. Through Markov-Chain Monte Carlo (MCMC) routines, BADASS performs simultaneous multi-component fits to emission-line spectra. Individual spaxels were fit iteratively across the full spatial extent of the data cubes, where a third order Legendre polynomial was used for the continuum. Emission lines were fit with either Gaussian (recombination, molecular, and fine-structure lines) or Lorentzian functions (Polycyclic Aromatic Hydrocarbon [PAH] features). A secondary Gaussian component to the fit was included based on the F-test of variances: F = , where is the standard deviation of the residuals using either a single or double Gaussian fit. If F 3.0, then adding a second component provided a significant improvement to the fit and was thus justifiable. This secondary component was restricted to have a larger velocity offset from the rest-frame wavelength than the main component. Due to the diverse emission profiles in our data, we left the amplitude and width as unrestricted free parameters. We also set a signal-to-noise (S/N) threshold of 5 for all lines and components to determine a detection. Uncertainties were derived from the error spectrum and the random errors associated with the Bayesian fitting process.
@@ 4.1 Near-Infrared Spectra
In Figure 1, we show images and NIRSpec spectra of the two nuclei and two nearby regions in the disks of NGC 3256. The spectra shown were extracted using an aperture of radius . In the nuclear and off-nuclear regions of both galaxies, we detect a wealth of hydrogen recombination, H, He i, and [Fe ii] emission lines.
@@
The spectrum of the northern nucleus is similar to that of the off-nuclear regions. They both show a flat continuum and weak HO (3.05 m) and CO (4.27 m) ice absorption. The PAH emission feature at 3.3 m is also clearly seen. In contrast, the S nucleus shows a rising MIR continuum and deep HO and CO ice absorption. The ice absorption in the S nucleus is significant, with the continuum level dropping by over an order of magnitude. There is also pronounced absorption in the CO rovibrational band at 4.67 m (Pereira-Santaella et al., 2024).
@@
Although the NIRSpec observations do not cover the full extent of the CO outflow regions as detected by ALMA (Sakamoto et al., 2014; Michiyama et al., 2018), the regions where there is overlap show blue and red asymmetric profiles in many molecular and ionized emission lines, including [Fe ii] 1.25 m, H 1.96 m, and H 2.12 m, that are not seen elsewhere. These blue and red emission wings are detected to within 100 pc of the S nucleus and extend beyond the NIRSpec FoV. In order to study the full extent of the H emission, much of our analysis will focus on the MIRI data, which cover a larger area than NIRSpec.
4.2 Warm H Gas
@@
The wavelength coverage of JWST/MIRI allows us to examine the warm H gas using the H 0–0 S(8)–S(1) emission lines from 5.05 to 17.0 m. The larger spatial coverage of MIRI (compared to NIRSpec) allows us to better trace the extent of the H emission. In Figure 2, we plot the full MIRI spectra extracted from the regions where we detect the strongest emission observed in all MIRI channels of the secondary, asymmetric H component. These locations are about 650 and 350 pc to the north and south of the S nucleus. The continuum and emission features of both spectra are similar, albeit the continuum level of the redshifted gas is 3–8 times higher than the blueshifted gas, though this could be due to the different aperture locations relative to the S nucleus.
@@
A sample of the diverse H line profiles in our data are shown in Figure 3, along with the best fit to the main and secondary components, where the main component is defined as the emission profile closest to the systemic velocity. We also plot the velocity maps of the main and secondary H S(1) components in central panels of Figure 3. In the following two subsections, we describe the flux distribution and kinematics of the main and secondary components.
4.2.1 Main H Gas Component
@@
The warm H gas structure of the main component in the northern galaxy appears largely undisturbed while the gas in the southern galaxy appears asymmetric and complex. Specifically, as shown in the first column, first and third rows of Figure 4, there is a horizontal feature, likely the southern galactic disk seen edge-on, that cuts across the S nucleus (English et al., 2003). There is also a vertical structure that reaches a peak luminosity roughly 350 pc south of the S nucleus. In terms of the kinematics, the redshifted emission seen in the northeast extends to the south, where it reaches a maximum velocity of +130 km s about 300 pc east of the S nucleus (near extraction region C in Figure 3). A similar structure of blueshifted gas to the west of the S nucleus is noticeably absent. Lastly, the velocity dispersion also shows a complex morphology throughout the FoV, and is largest immediately to the north and south of the southern nucleus in the same regions where the secondary component is detected. In contrast, it is comparatively low in the regions surrounding the N nucleus.
4.2.2 Secondary H Gas Component

### v3

@@ Abstract
We present James Webb Space Telescope (JWST) Integral Field Spectrograph observations of NGC 3256, a local infrared-luminous late-stage merging system with two nuclei roughly 1 apart, both of which have evidence of cold molecular outflows. Using JWST/NIRSpec and MIRI datasets, we investigate this morphologically complex system on spatial scales of 100, where we focus on the warm molecular H2 gas surrounding the nuclei. We detect collimated outflowing warm H2 gas originating from the southern nucleus, though we do not find significant outflowing H2 gas surrounding the northern nucleus. We measure maximum intrinsic outflow velocities of 1,000, which extend out to a distance of 0.7. Based on H2 S(7)/S(1) ratios, we find a larger fraction of warmer gas near the S nucleus, which decreases with increasing distance from the nucleus, signifying the southern nucleus as a primary source of H2 heating. The gas mass of the warm H2 outflow component is estimated to be , as much as 6 of the cold H2 mass estimated using ALMA CO data. The outflow time scale is about , resulting in a mass outflow rate and kinetic power . Lastly, regions within our NIRSpec data where the outflowing gas reside show high [Fe II]/Pa and H2/Br line ratios, indicate enhanced mechanical heating caused by the outflows. The fluxes and ratios of Polycyclic Aromatic Hydrocarbons (PAH) in these regions are not significantly different compared to those elsewhere in the disk, suggesting the outflows may not significantly alter the PAH ionization state or grain size.
@@ 1 Introduction
The physical processes that govern the evolution of galaxies are of fundamental importance to our understanding of how galaxies in the Universe formed. These processes manifest in many forms, including internal mechanisms, such as stellar and active galactic nucleus (AGN) outflows (e.g., Müller-Sánchez et al., 2011; U et al., 2019; Aravindan et al., 2023), and external processes such as galaxy mergers (e.g., Lin et al., 2008; Privon et al., 2013; Conselice, 2014). Outflows are believed to regulate and suppress star formation and can, in many instances, lead their hosts to the well-defined red sequence (e.g., Croton et al., 2006; King & Pounds, 2015; Veilleux et al., 2020). Major mergers also play a critical role in the formation of massive galaxies, and are attributed to the coevolution of the host galaxy and supermassive black holes (e.g., Medling et al., 2015; Ricci et al., 2017; Ramos Almeida et al., 2019) and the formation of compact star clusters (e.g., Mulia et al., 2015; Adamo et al., 2020; Linden et al., 2021). Local luminous infrared galaxies (LIRGs; ) are ideal systems for studying these processes because they allow us to analyze evolutionary processes across the full array of merger states, and evaluate the elevated star formation rates triggered by such mergers. Their diversity in starburst and AGN power, merger stages, and outflow activity makes LIRGs compelling ecosystems for tying these properties to the stages of galaxy evolution.
@@
One of the most compelling aspects of NGC 3256 is the presence of outflows in both nuclei. Submillimeter ALMA observations indicate two different biconical outflows of molecular CO gas originating from the nuclei, with deprojected maximum velocities (and inclination angles) of 750 (60∘) and 2600 km s-1 (20∘) for the outflows in the northern and southern galaxies, respectively (Sakamoto et al., 2014; Michiyama et al., 2018). Near-infrared (NIR) VLT/SINFONI observations of the S nucleus also reveal biconical outflows in the warm, 2000 K H2 gas, where the gas mass of these outflowing components is estimated to be 1,200 (Emonts et al., 2014).
While previous observations have been able to analyze the warm and cold molecular gas phases in NIR and submillimeter wavelengths with great detail, analysis of the warm H2 component in the mid-infrared (MIR) has been mainly restricted to using the S(1)–S(3) transitions due to the resolution and sensitivity of the Spitzer Space Telescope (e.g. Stierwalt et al., 2013; Inami et al., 2013). Petric et al. (2018) calculated the mass of the warm molecular H2 component of the northern galaxy to be log() = 7.38 and a temperature of 322 K within the IRS slits that were centered on the N nucleus. While Spitzer gave a great first look at the warm H2 gas, the H2 emission was not spatially resolved and the spectral resolution of Spitzer, R60–600, made it difficult to spectrally resolve the outflowing gas in NGC 3256.
The various H2 transitions and their line ratios with other IR lines are useful for identifying the sources heating the molecular gas and energizing the interstellar medium (ISM; Larkin et al., 1998; Riffel et al., 2013; Colina et al., 2015). They can be used to identify the presence of shocks induced by the outflows, and help evaluate the role outflows have in regulating star formation. With JWST, we can now perform the most detailed, spatially resolved analysis of outflowing warm H2 gas to date in NGC 3256. The unprecedented spatial resolution of JWST enables us to examine the warm H2 component on scales of 40–100 pc. We also take advantage of the superb sensitivity and wavelength coverage of JWST to analyze the H2 0–0 S(1)–S(8) rotational lines between 5–17 m. Specifically, we utilize integral-field spectroscopy (IFS) of the medium-resolution spectrometer (MRS) of the Mid-Infrared Instrument (MIRI, Rieke et al., 2015; Labiano et al., 2021; Wright et al., 2023) and the Near-InfraRed Spectrograph (NIRSpec, Böker et al., 2022; Jakobsen et al., 2022) to assess the kinematics and energetics of the outflows, and evaluate their impact on the local ISM.
In our analysis, we adopt a cosmology of km s-1 Mpc-1, , and . At the redshift of NGC 3256, (40.4 Mpc), subtends 190 pc.
@@ 2 Observations and Data Reduction
Observations of NGC 3256 were taken as part of the Early Release Science (ERS) Program 1328 (co-PIs: L. Armus and A. Evans), for which the JWST/NIRSpec IFS data were taken on 2022 December 23-24. Two sets of observations, each centered on a nucleus, were taken using three high-resolution gratings: G140H/F100LP, G235H/F170LP, and G395H/F290LP. This resulted in a wavelength coverage of 0.97–5.27 m, with a nominal resolving power of R2700. A four-point dither pattern was utilized, and the total exposure times for all gratings were 934 s and 2218 s for the northern and southern pointings, respectively. Observations using JWST/MIRI in the MRS mode were taken on 2022 December 27. Like for NIRSpec, a separate pointing for each nucleus was taken. These observations include all three subbands (short [A], medium [B], and long [C]) in all four channels, which cover the full wavelength range from 4.9–28.8 m. The total exposure time for each subband/channel was 444 s, and was the same for both pointings. In this article, we predominately use channels 1–3, which have a nominal resolving power of R2000–3700.
The data were processed through the standard JWST Calibration Pipeline (ver.1.12.5 for NIRSpec; ver.1.11.3 for MIRI, Bushouse et al., 2023). Stage 1 of the reduction pipeline, Detector1, implements detector-level corrections and outputs count-rate data products. Stage 2 (spec2) applies various calibrations steps, including distortion, wavelength, and flux calibrations. Stage 3 (spec3) performs background subtraction for MIRI data and builds the final data cubes from the individual exposures. We also incorporated leakcal calibration in our NIRSpec data. Before stage 3 was implemented, we only included spaxels with data quality (DQ) equal to 0 in order to remove bad spaxels contaminating the NIRSpec data products. We additionally flagged any spaxels in the postpipeline products with values significantly above emission lines or that were negative.
To properly align both the NIRSpec and MIRI data, Gaia DR3 (Gaia Collaboration et al., 2016, 2023) reference frames of nearby field stars were first used to derive WCS corrections for the JWST/NIRCam and MIRI images (also obtained by ERS 1328, see Linden et al., 2024). The IFS cubes were subsequently collapsed into a single frame and aligned to the WCS corrected NIRCam/MIRI images, where corrections were R.A.: + and decl.: + for NIRSpec, and R.A.: and decl.: + for MIRI.
Lastly, fringing corrections were applied to all extracted 1D MIRI spectra using the calibration pipeline. We also stitched the spectra by scaling all the continua to that of MIRI channel 3 long, the longest wavelength channel used in our analysis, where scaling factors were typically around 5.
@@ 3 Methods of Analysis
The primary focus of this article is to obtain robust mass estimates and temperatures of the warm outflowing H2 gas, as well as analyze its kinematics and energetics. Due to the complex kinematic nature of this merging system, robust emission-line decompositions and detailed mapping of the spatial structures are needed to accurately identify and study the H2 gas. To accomplish this, we used the Bayesian AGN Decomposition Analysis for SDSS Spectra (BADASS111https://github.com/remingtonsexton/BADASS3, Sexton et al., 2021) code. Through Markov-Chain Monte Carlo (MCMC) routines, BADASS performs simultaneous multicomponent fits to emission-line spectra. Individual spaxels in both the NIRSpec and MIRI datasets were fit iteratively across the full spatial extent of the data cubes, where a third-order Legendre polynomial was used for the continuum. Emission lines were fit with either Gaussian (recombination, molecular, and fine-structure lines) or Lorentzian functions (polycyclic aromatic hydrocarbon (PAH) features). A secondary Gaussian component to the fit was included based on the F-test of variances: F = , where is the standard deviation of the residuals using either a single or double Gaussian fit. If F 3.0, then adding a second component provided a significant improvement to the fit and was thus justifiable. This secondary component was restricted to have a larger velocity offset from the rest-frame wavelength than the main component. Due to the diverse emission profiles in our data, we left the amplitude and width as unrestricted free parameters. We also set a signal-to-noise ratio (S/N) threshold of five for all lines and components to determine a detection. Uncertainties were derived from the error spectrum and the random errors associated with the Bayesian fitting process.
@@ 4.1 Near-infrared Spectra
In Figure 1, we show NIRCam images and NIRSpec spectra of the two nuclei and two nearby regions in the disks of NGC 3256. The spectra shown were extracted using an aperture of radius . In the nuclear and off-nuclear regions of both galaxies, we detect a wealth of hydrogen recombination, H2, He I, and [Fe II] emission lines.
@@
The spectrum of the N nucleus is similar to that of the off-nuclear regions. They both show a flat continuum and weak H2O (3.05 m) and CO2 (4.27 m) ice absorption. The PAH emission feature at 3.3 m is also clearly seen. In contrast, the S nucleus shows a rising MIR continuum and deep H2O and CO2 ice absorption. The ice absorption in the S nucleus is significant, with the continuum level dropping by over an order of magnitude. There is also pronounced absorption in the CO rovibrational band at 4.67 m (Pereira-Santaella et al., 2024).
@@
Although the NIRSpec observations do not cover the full extent of the CO outflow regions as detected by ALMA (Sakamoto et al., 2014; Michiyama et al., 2018), the regions where there is overlap show blue and red asymmetric profiles in many molecular and ionized emission lines, including [Fe II] 1.25 m, H2 1–0 S(3) 1.96 m, and H2 1–0 S(1) 2.12 m, that are not seen elsewhere. These blue and red emission wings are detected to within 100 pc of the S nucleus and extend beyond the NIRSpec FOV. In order to study the full extent of the H2 emission, much of our analysis will focus on the MIRI data, which cover a larger area than NIRSpec.
4.2 Warm H2 Gas
@@
The wavelength coverage of JWST/MIRI allows us to examine the warm H2 gas using the H2 0–0 S(8)–S(1) emission lines from 5.05 to 17.0 m. The larger spatial coverage of MIRI (compared to NIRSpec) allows us to better trace the extent of the H2 emission. In Figure 2, we plot the full MIRI spectra extracted from three regions where we detect emission of the secondary, asymmetric H2 component observed in all MIRI channels. One aperture was placed 350 pc south of the S nucleus where we measure strong blueshifted gas emission, while two other apertures were placed 350 and 650 pc north of the S nucleus where we detect redshifted gas emission. The continuum and emission features of all spectra are similar, albeit the continuum level of the redshifted gas is as much as 3 times higher than the blueshifted gas for regions equidistant from the S nucleus.
@@
A sample of the diverse H2 S(1) line profiles in our data are shown in Figure 3, along with the best fit to the main and secondary components, where the main component is defined as the emission profile closest to the systemic velocity. We also plot the velocity maps of the main and secondary H2 S(1) components in central panels of Figure 3. In the following two subsections, we describe the flux distribution and kinematics of the main and secondary components.
4.2.1 Main H2 Gas Component
@@
The warm H2 gas structure of the main component in the northern galaxy appears largely undisturbed while the gas in the southern galaxy appears asymmetric and complex. Specifically, as shown in the first column, first and third rows of Figure 4, there is a horizontal feature, likely the southern galactic disk seen edge on, that cuts across the S nucleus (English et al., 2003). There is also a vertical structure that reaches a peak luminosity roughly 350 pc south of the S nucleus. In terms of the kinematics, the redshifted emission seen in the northeast extends to the south, where it reaches a maximum velocity of +130 km s-1 about 300 pc east of the S nucleus (near extraction region C in Figure 3). A similar structure of blueshifted gas to the west of the S nucleus is noticeably absent. Lastly, the velocity dispersion also shows a complex morphology throughout the FOV, and is largest immediately to the north and south of the S nucleus in the same regions where the secondary component is detected. In contrast, it is comparatively low in the regions surrounding the N nucleus.
4.2.2 Secondary H2 Gas Component